a rectangular has width of 1800 cm and length of 2400 cm. the floor will be covered with tiles. each tile . is a square of side 30 cm calculate the number of tiles required to cover the floor fully.
1
Expert's answer
2011-10-21T10:02:56-0400
Floor area is Sf = 1800*2400 = 4320000 cm^2 and the area of a tile is St = 30^2 = 90 cm^2. So, we need N = Sf/St = 4320000/90 = 48000 tiles.